How much energy would have been required to warm one mole of liquid water from 0 C to its boiling point?
1
Expert's answer
2020-01-31T07:53:24-0500

"Q = cm\Delta T"

"C(water) = 4.184 \frac{J}{g \times ^\circ C}"

Find mass of 1 mole of liqiud water

"m= M\times n = 18.02\frac{g}{mol}\times 1 mol = 18.02 g"


"Q = 4.184\times18.02 \times (100 ^\circ C - 0 ^\circ C)=7539.6 J"